OpenStudy (anonymous):

find the quadratic polynomial whose graph goes through the points (-2,5),(0,1),(2,13)

OpenStudy (campbell_st):

plot the points  so you know its concave up the y intercept is 1 so its in the form \[y = ax^2 + bx + 1\] so substituting x = -2 y is 5 so 4a -2b + 1 = 5 Equation 1 and substituting x = 2 4a + 2b + 1 = 13 Equation 2 solve equations 1 and 2 simultaneously by adding the equations to eliminate b 4a - 2b + 1 = 5 4a + 2b + 1 = 13 --------------- 8a + 2 = 18 then a = 2 substitute a into equation 1 to find b then you'll have your quadratic...